OpenStudy (anonymous):

Annabelle and Navene are multiplying (4275)(4372). Annabelle's Work Navene's Work (4275)(4372) = 42 + 375 + 2 = 4577 (4275)(4372) = 42⋅375⋅2 = 46710 Is either of them correct? Explain your reasoning.

OpenStudy (anonymous):

I think the two answers got a little messed up...could you maybe take a screenshot of this problem?

OpenStudy (anonymous):

Annabelle and Navene are multiplying (4^2 7^5)(4^3 7^2). Annabelle's Work (4^2 7^5)(4^3 7^2) = 42 + 375 + 2 = 4^5 7^7 Navene's Work (4^2 7^5)(4^3 7^2) = 42⋅375⋅2 = 4^6 7^10 Is either of them correct? Explain your reasoning.

OpenStudy (anonymous):

This is actually a little beyond me. Sorry. I'm sure someone else can help you out. I'd rather have someone give you a good answer than I try to give you an answer and totally fail. :)

OpenStudy (anonymous):

its cool i understand

OpenStudy (triciaal):

is it possible to see the original question? (4^2 7^5)(4^3 7^2) you seem to be missing operation sign do you mean (4^2*7^5)(4^3*7^2)? in which case your 4577 is correct

OpenStudy (triciaal):